OpenStudy (anonymous):

what does 7 multiplied by 5 multiplied by 2 whole over 7 multiplied by 3, the whole raised to the power of 2 multiplied by 5 to the power of 0 over 2 to the power of negative 3, whole to the power of 3 multiplied by 2 to the power of negative 9 equal to?

OpenStudy (mathteacher1729):

Can you type that out using math notation? I assume the first part of the problem means this: > 7 multiplied by 5 multiplied by 2 whole over 7 multiplied by 3, the whole raised to the power of 2 \[ \left(\dfrac{7\cdot 5\cdot 2}{7 \cdot 3}\right)^2 \] But the way you have it written there is a great deal of ambiguity to the order of operations and what I have written above could be incorrect.

OpenStudy (anonymous):

NO what you wrote is correct

OpenStudy (mathteacher1729):

Ok, what about this? > raised to the power of 2 multiplied by 5 to the power of 0 over 2 to the power of negative 3, whole to the power of 3 multiplied by 2 to the power of negative 9 equal to? Can you type that in mathematical notation? (Using correct parenthesis and/or \(\LaTeX\)?)

OpenStudy (quickstudent):

Welcome to Openstudy, @19marrow ! Solving this equation is actually very easy. You just need to follow a couple steps for it.

OpenStudy (quickstudent):

When working with any equation with operations in the numerator or denominator, you have to solve for the numerator first. In other words, multiply 7 * 5 * 2. What does that equal?

OpenStudy (quickstudent):

Next, you solve for the denominator. That means multiplying 7 * 3. What is that?

OpenStudy (quickstudent):

Find the answers to these and then I will tell you the next steps.